OpenStudy (anonymous):

Help ASAP For the inverse variation equation xy = k, what is the constant of variation, k, when x = –3 and y = –2

OpenStudy (anonymous):

–6 -2/3 3/2 6

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (anonymous):

-2/3 would be my guess

OpenStudy (skullpatrol):

Guessing randomly doesn't work in math :-) Please give a reason.

OpenStudy (anonymous):

because usually its x/y right? so the inverse would be y/x

OpenStudy (skullpatrol):

You are correct that "usually" you see expressions like x/y in class and the inverse of x/y is y/x, but that is not what this question is asking for.

OpenStudy (anonymous):

So its 6?

OpenStudy (anonymous):

no that doesnt make sense

OpenStudy (skullpatrol):

Yes. 6 is correct :-)

OpenStudy (skullpatrol):

For the inverse variation equation xy = k, what is the constant of variation, k, when x = –3 and y = –2 x * y = k -3 (-2) = 6

OpenStudy (anonymous):

Thank you! Can you help me with another one?

OpenStudy (skullpatrol):

I can try.

OpenStudy (anonymous):

The inverse variation equation shows the relationship between wavelength in meters, x, and frequency, y. y=3x10^8/x What is the wavelength for radio waves with frequency 3 × 10^9? 1 × 10^–1 m 3 × 10^–1 m 3 × 10^17 m 9 × 10^17 m

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (skullpatrol):



OpenStudy (anonymous):

so find y=3x10^8/x

OpenStudy (skullpatrol):

Read the problem carefully. Decide what unknown number is being asked for.

OpenStudy (anonymous):

3x10^9?

OpenStudy (skullpatrol):

that's not an answer choice

OpenStudy (anonymous):

I still dont understand

OpenStudy (skullpatrol):

what is the question asking for?

OpenStudy (anonymous):

What is the wavelength for radio waves with frequency 3 × 10^9

OpenStudy (skullpatrol):

Good, where is the wavelength in the formula they gave you?

OpenStudy (anonymous):

wavelength is x and frequency is y

OpenStudy (skullpatrol):

so they want you to find x, right?

OpenStudy (anonymous):

yes

OpenStudy (skullpatrol):